Direct Materials Variances

The following data relate to the direct materials cost for the production of 50,000 automobile tires:

Actual: 725,000 lbs. at $3.00 per lb.
Standard: 730,000 lbs. at $2.95 per lb.

a. Determine the direct materials price variance, direct materials quantity variance, and total direct materials cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct Materials Price Variance $
Direct Materials Quantity Variance $
Total Direct Materials Cost Variance $

b. The direct materials price variance should normally be reported to the  . When lower amounts of direct materials are used because of production efficiencies, the variance would be reported to the  . When the favorable use of raw materials is caused by the purchase of higher-quality raw materials, the variance should be reported to the  .
Answer #1
a
Direct Materials Price Variance 36250 Unfavorable =725000*(3-2.95)
Direct Materials Quantity Variance -14750 Unfavorable =2.95*(725000-730000)
Total Direct Materials Cost Variance 21500 Unfavorable =(725000*3)-(730000*2.95)
b
The direct materials price variance should normally be reported to the Purchasing Department.When lower amounts of direct materials are used because of production efficiencies, the variance would be reported to the production supervisor. When the favorable use of raw materials is caused by the purchase of higher-quality raw materials, the variance should be reported to the Purchasing Department.
Answer #2
Negative numbers are favorable.